# Plasticine: Accelerating Research in Plasticity-Motivated Deep Reinforcement Learning

### v1

@@ Appendix G Experiments
Working in progress…
G.1 Standard RL Scenario
G.1.1 PPO+Procgen
Table 2: Default hyperparameters used for the PPO+Procgen experiments.
@@ G.1.1 PPO+Procgen
Observation downsampling (64, 64) Optimizer Adam
Observation normalization / 255. Learning rate 5e-4
Reward normalization No GAE coefficient 0.95
Weight initialization Orthogonal Action entropy coefficient 0.01
LSTM No Value loss coefficient 0.5
Stacked frames No Value clip range 0.2
Environment steps 25000000 Max gradient norm 0.5
Episode steps 256 Epochs per rollout 3
Number of workers 1 Batch size 2048
Environments per worker 64 Discount factor 0.999
G.1.2 PQN+ALE
Table 3: Default hyperparameters used for the PQN+ALE experiments.
@@ G.1.2 PQN+ALE
Observation downsampling (84, 84) Starting epsilon for exploration 1
Observation normalization / 255. Ending epsilon for exploration 0.01
Reward normalization No Fraction of exploration 0.1
Weight initialization Orthogonal Lambda for Q-learning 0.65
LSTM No GAE coefficient 0.95
Stacked frames 4 Action entropy coefficient 0.01
Environment steps 10000000 Value loss coefficient 0.5
Episode steps 128 Value clip range 0.2
Number of workers 1 Max gradient norm 10.0
Environments per worker 8 Epochs per rollout 4
Optimizer Adam Batch size 256
Learning rate 2.5e-4 Discount factor 0.99
G.1.3 TD3+DMC
Table 4: Default hyperparameters used for the TD3+DMC experiments.
@@ G.1.3 TD3+DMC
Warm-up steps 1e4 Environments per worker 1
Replay buffer size 1e6 Optimizer Adam
State normalization No Learning rate 1e-4
Reward normalization No Target smoothing coefficient 0.005
Weight initialization Orthogonal Target policy noise 0.2
LSTM No Target noise clip 0.5
Non-linearity ReLU Policy delay 2
Environment steps 5e6 Batch size 256
Number of workers 1 Discount factor 0.99
G.2 Continual RL Scenario
G.3 Open-ended RL Scenario
Table 5: Default hyperparameters used for the PPO+Craftax experiments.
Hyperparameter Value Hyperparameter Value
Observation downsampling N/A Optimizer Adam
Observation normalization None Learning rate 2e-4
Reward normalization No GAE coefficient 0.8
Weight initialization Orthogonal Action entropy coefficient 0.01
LSTM No Value loss coefficient 0.5
Stacked frames N/A Value clip range 0.2
Environment steps 100M Max gradient norm 1.0
Episode steps 64 Epochs per rollout 4
Number of workers 1 Batch size 8192
Environments per worker 1024 Discount factor 0.99
Appendix H Author Contributions
We summarize the main contributions from each author as follows (names are listed in alphabetical order):
  • Project Lead: Mingqi Yuan
  • Algorithm Development: Mingqi Yuan
  • Environment Development: Qi Wang
  • Metric Development and Project Evaluation: Guozheng Ma
  • Area Advisor: Bo Li, Tao Yu, Xin Jin, and Yunbo Wang
  • Supervision: Dacheng Tao, Wenjun Zeng, and Xiaokang Yang

### v2

@@ Appendix G Experiments
In this section, we conduct benchmark experiments for all three learning scenarios and compare the performance of the implemented approaches. The details of these experimental environments can be found in Appendix E. The following figures only show a part of the experiment results, and more detailed benchmark experiment results can be found in Plasticine’s Wandb space [^1] https://wandb.ai/yuanmingqi/Plasticine/reportlist.
G.1 Plasticine+C51+ALE
@@
Number of workers 1 Number of atoms ($N$) 51
Environments per worker 1 Support lower bound ($V_{min}$) -10
Total timesteps 1e7 Support upper bound ($V_{max}$) 10
Learning rate 2.5e-4 Replay buffer size 1e6
Discount factor 0.99 Batch size 32
Target network frequency 10000 Start epsilon ($\epsilon$) 1
Learning starts 80000 End epsilon ($\epsilon$) 0.01
Train frequency 4 Exploration fraction 0.10
Table 1: Hyperparameters used for the C51 agent in the standard ALE experiments.
Figure 5: Performance comparison of the vanilla C51 agent and its combinations with several Plasticine methods on the ALE - DemonAttack environment. The solid line and shaded region represent the mean and standard deviation across multiple runs, respectively.
G.2 Plasticine+PPO+Continual Procgen
@@
Number of tasks 10 Level offset 20
Observation downsampling (84, 84) Optimizer Adam (default)
Observation normalization / 255. Learning rate 1e-3
Reward normalization Yes GAE coefficient 0.95
LSTM No Entropy coefficient 0.01
Stacked frames 4 Value loss coefficient 0.5
Environment steps (per round) 2e6 Value clip range 0.2
Episode steps 1000 Max gradient norm 0.5
Number of workers 1 Number of mini-batches 8
Environments per worker 1 Discount factor 0.99
Table 2: Hyperparameters used for the PPO agent in the continual procgen experiments.
Figure 6: Performance comparison of the vanilla PPO agent and its combinations with several Plasticine methods on the Continual Procgen - StarPilot environment. The solid line and shaded region represent the mean and standard deviation across multiple runs, respectively.
G.3 Plasticine+PPO+Continual DMC
@@
Learning rate 3e-4 Optimizer Adam (default)
Reward normalization Yes GAE coefficient 0.95
LSTM No Entropy coefficient 0.0
Initial std. 1.0 Value loss coefficient 0.5
Environment steps (per round) 1e6 Value clip range 0.2
Episode steps 2048 Max gradient norm 0.5
Number of workers 1 Number of mini-batches 32
Environments per worker 1 Discount factor 0.99
Table 3: Hyperparameters used for the PPO agent in the continual DMC experiments.
Figure 7: Performance comparison of the vanilla PPO agent and its combinations with several Plasticine methods on the Continual DMC - Walker environment. The solid line and shaded region represent the mean and standard deviation across multiple runs, respectively.
Acknowledgements
This work is supported, in part, by the Hong Kong SAR Research Grants Council under Grant No. PolyU 15224823, the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2024A1515011524, the NSFC under Grant No. 62302246, the ZJNSFC under Grant No. LQ23F010008, the Ningbo under Grants No. 2023Z237 & 2023CX050011 & 2024Z284 & 2024Z289 & 2025Z038 & 2025Z059, and the Ningbo Institute of Digital Twin (IDT) under Grant No. S203.2.01.32.002. We thank the high-performance computing center at INFIFORCE Intelligent Technology Co., Ltd., Eastern Institute of Technology, and IDT for providing the computing resources.